Company A uses the periodic inventory system. The beginning inventory consisted of 360 units that cost $65 each. During the month of febuary, the company made two purchases: on febuary 15, they bought 450 units at $68 each; on febuary 25, they bought 270 units at $70 each. Company A also sold 900 units during the month

Using LIFO, what is the amount of ending inventory.
Answer #1

Units available for sale = 360+450+270 = 1,080 units Units sold = 900 Units in ending inventory = 1,080 - 900 = 180 units Amo